Last year, Jodi invested $10,000, part at 6% annual interest and the rest at 8% annual interest. If she received $760 in interest at the end of the year, how much did she invest at each rate?
Answer by checkley77(12844)

.08x+.06(10,000-x)=760
.08x+600-.06x=760
.02x=760-600
.02x=160
x=160/.02
x=8,000 amount invested @ 8%
10,000-8,000=2,000 invested @ 6%
Proof:
.08*8,000+.06*2,000=760
640+120=760
760=760